OpenStudy (linna):

A product's price is 150$ . And it raises for 8% .... after a couple of days it gets down for 8% again. What's the new price? I keep getting the same answer. the answer is: 149.04$ (ATTEMPT SHOWN)

OpenStudy (linna):

\[\begin{array}{l}150(1 + 8\% ) = x\\150 + \frac{8}{{100}} = x\\15000 + 8 = 100x\\15008 = 100x\\150.08 = x\\\\150.08(1 - 8\% ) = x\\150.08 - \frac{8}{{100}} = x\\\frac{{150.08 \cdot 100 - 8}}{{100}} = x\\\frac{{15008 - 8}}{{100}} = x\\\frac{{15000}}{{100}} = x\\x = 150\end{array}\]

OpenStudy (linna):

leave it.. I know ahh... 150(1+8/100) is not 150+8/100